# Factorizable joint shift revisited

### v1

@@ Abstract
Factorizable joint shift (FJS) was proposed as a type of distribution shift (or dataset shift) that comprises both covariate and label shift. Recently, it has been observed that FJS actually arises from consecutive label and covariate (or vice versa) shifts. Research into FJS so far has been confined to the case of categorical label spaces. We propose a framework for analysing distribution shift in the case of general label spaces, thus covering both classification and regression models. Based on the framework, we generalise existing results on FJS to general label spaces and propose a related extension of the expectation maximisation (EM) algorithm for class prior probabilities. We also take a fresh look at generalized label shift (GLS) in the case of general label spaces.
Keywords: Distribution shift, dataset shift, factorizable joint shift, generalized label shift, domain adaptation.
@@ 1 Introduction
Classification and regression models may work very well on their respective training datasets but fail utterly when being deployed in test datasets. Such failure can be caused by distribution shift [^1] Cf. Zhang et al. (2024). Distribution shift is also known as dataset shift (Quiñonero-Candela et al., 2009; Moreno-Torres et al., 2012). between the training and test datasets. For this reason, distribution shift and domain adaptation (a notion comprising techniques for tackling distribution shift) has been a major research topic in machine learning for some time. This paper takes the perspective of Kouw and Loog (2021) and studies the case where feature observations from the test dataset are available for analysis but observations of labels are missing.
Under these circumstances, without any assumptions on the nature of the distribution shift between the training and the test datasets meaningful prediction of the labels in the test dataset or of their distribution is not feasible. See Kouw and Loog (2021) for a survey of approaches to domain adaptation and their related assumptions. Arguably, covariate shift (also known as population drift) and label shift (also known as prior probability shift or target shift) are the most popular specific distribution shift assumptions, both for their intuiveness as well as their computational manageability. However, exclusive covariate and label shift assumptions have been criticised for being insufficient for common domain adaptation tasks (e.g. Tachet des Combes et al., 2020).
He et al. (2021) proposed the notion of factorizable joint shift (FJS) as an assumption for distribution shift that covers both covariate and label shift. Recently, Dong et al. (2025) observed that FJS actually may be described as combination of label and covariate shifts, thus making the assumption more plausible. Tasche (2022) analysed FJS for categorical label spaces with regard to the potential and limitations of domain adaptation based on this assumption. This paper updates the analyses of Tasche (2022) for FJS in the case of general label spaces. More precisely, the main contributions of this paper to the subject of distribution shift are the following:
  • Pointing out the building blocks of non-specific distribution shift with general label spaces (including Euclidean, categorical and mixed Euclidean-categorical spaces) in order to facilitate the demarcation and analysis of specific types of distribution shift.
@@
  • Proposing a generalisation to general label spaces of the expectation maximisation (EM) algorithm for estimating target prior class probabilities by Saerens et al. (2002) .
@@
The setting and notation for this paper are set out in Section 2, followed by a framework of general assumptions and results on distribution shift. Section 3 comprises analyses and results regarding FJS as well as approaches to estimating the characterics of FJS, including the proposal for a general EM algorithm. The relation between GLS and FJS is analysed in Section 4. Section 5 concludes the paper, with a summary and suggestions for further research. Appendix A presents a description of the classification approach to density estimation.
@@ 2 The building blocks of distribution shift
We study distribution shift related to classification and regression problems in a probabilistic setting with feature (also known as input, explanatory or independent) variables $X$ and label (also known as output, response or dependent) variables $Y$. We assume that a joint source (also known as training) probability distribution $P=P_{(X,Y)}$ of $X$ and $Y$ already has been learnt. Now we are presented with a joint target (also known as test) distribution $Q=Q_{(X,Y)}$ of $X$ and $Y$ which is only partially observable. The problem is to infer information about the posterior (also known as conditional) distribution $Q_{Y|X}$ of $Y$ given $X$ under the target distribution $Q$ from the source distribution $P$ despite a lack of joint observations of $(X,Y)$ under $Q$. Lack of observations from the target distribution $Q$ and more generally lack of information about $Q$ becomes an issue in the presence of distribution shift, i.e. if the source and target distributions possibly differ (Zhang et al., 2024, Section 4.9). In Section 3 below, we analyse the two cases where
@@
The problem phrased in (B) is an unsupervised domain adaptation problem where a classification or regression function is learnt on a completely known dataset in a supervised manner and is then adapted to another dataset without label information in an unsupervised manner (Kouw and Loog, 2021, Section 6.7). Problem (A) where only joint observations of features and label are missing has less frequently been looked at in the machine learning literature. It is akin to copula estimation (Durante and Sempi, 2016) and as such has some practical relevance (see e.g. Jaworski et al., 2013; Tasche, 2025). Problem (A) is investigated in this paper because its solutions can inform approaches to solutions for Problem (B).
@@
  • knowledge of the building blocks can inform generative approaches to distribution shift when datasets are simulated.
@@ 2.1 Setting
The setting for this paper is a standard setting for domain adaptation and therefore, with the exceptions of the notation and the fact that the label and feature spaces considered are more general, similar to the setting described for instance in Section 1.1 of Scott (2019).
@@
  • The source distribution $P$ and the target distribution $Q$ are different, i.e. $P\neq Q$. In other words, there is a distribution shift (see e.g. Zhang et al., 2024) between source and target. Occasionally, non-specific distribution shift also is called dataset shift in the literature (Storkey, 2009; Moreno-Torres et al., 2012).
@@ 2.2 Notation
Expected values and probabilities conditional on $\sigma$-algebras or on random variables are well-defined for all combinations of probability spaces and integrable or non-negative random variables subject to the conditioning (Bauer, 1981, Sections 10.1 and 10.2; Klenke, 2013, Sections 8.2 and 8.3).
@@
In general, the existence of regular conditional distributions can be guaranteed only if certain conditions on the conditioned measurable space – $(\Omega_{Y},\mathcal{G})$ in Definition 2.1 – are satisfied (Bauer, 1981, Section 10.3; Klenke, 2013, Section 8.3). If however the label variable $Y$ is real-valued, the regular conditional distribution $P_{Y|X}$ exists, without any restrictions on the feature variable $X$ (Klenke, 2013, Theorem 8.28).
@@ 2.3 Linking source and target distributions in the face of distribution shift
In the literature the ‘transfer’ of information from $P$ to $Q$ by means of (2.5b) is sometimes called ‘importance reweighting’ and the density $f$ is referred to as ‘importance weights’ (e.g. Zhang et al., 2013; Kouw and Loog, 2021). More often than not, in the literature the reference measure for the density $f$ remains undisclosed or appears to be the Lebesgue measure (e.g. Zhang et al., 2013; Lipton et al., 2018). A notable exception is Scott (2019) who however proposes a reference measure $\mu$ that is available only in the case of a categorical label space.
@@
As application of Fubini’s theorem (2.4a) cannot be justified in the absence of any assumptions on the properties of the source distribution $P$ in the setting specified in Section 2.1, it is not possible to draw further conclusions about the composition of the density $f$ or the relationship between the posterior probabilities of $X$ given $Y$ under the target distribution $Q$ and the source distribution $P$ respectively. This is in contrast to the situation studied by Tasche (2022). In Tasche (2022), thanks to the assumption that the label variable is categorical, more refined formulae for $f$ (Theorem 1) and the posterior probabilities (Corollary 2, a more general version of Eq. (2.4) of Saerens et al., 2002, but of similar shape) could be derived.
@@ 2.4 Assumptions on the source distribution and conditional densities
    Under Assumptions 2.2 and 2.4 (i), let $h$ be a density of $Q_{X}$ with respect to $P_{X}$ and define $q_{Y|X}:\Omega_{X}\times\Omega_{Y}\to[0,\infty)$ by $q_{Y|X=x}(y)=\begin{cases}\frac{f(x,y)}{h(x)},&h(x)>0\\ 0,&h(x)=0\end{cases}$ for $x\in\Omega$ and $y\in\Omega_{Y}$. Then for $x\in\Omega$ with $h(x)>0$ (and therefore $Q_{X}$-a.s.) $q_{Y|X=x}$ is a density of $Q_{Y|X=x}$ with respect to $P_{Y|X=x}$.
@@
    Under Assumptions 2.2 and 2.4 (ii), let $g$ be a density of $Q_{Y}$ with respect to $P_{Y}$ and define $q_{X|Y}:\Omega_{X}\times\Omega_{Y}\to[0,\infty)$ by $q_{X|Y=y}(x)=\begin{cases}\frac{f(x,y)}{g(y)},&g(y)>0\\ 0,&g(y)=0\end{cases}$ for $x\in\Omega$ and $y\in\Omega_{Y}$. Then for $y\in\Omega$ with $g(y)>0$ (and therefore $Q_{Y}$-a.s.) $q_{X|Y=y}$ is a density of $Q_{X|Y=y}$ with respect to $P_{X|Y=y}$.
@@
with a fixed arbitrary $y_{0}\in\Omega_{Y}$, for $x\in\Omega_{X}$ and $G\in\mathcal{G}$. Then $K$ satisfies (i) of Definition 2.1 and is a measure on $(\Omega_{Y},\mathcal{G})$. By Fubini’s theorem (2.4a), it follows for $H\in\mathcal{H}$ and $G\in\mathcal{G}$ that
@@
(i) is basically the definition of $q_{X|Y=y}(x)$ in Thorem 2.6 (ii). (ii) is the combination of (i) and Corollary 2.5 (i). (iii) is implied by (i) and Corollary 2.5 (i). ∎
@@ 2.5 Facilitating calculations by further assumptions
Under Assumption 2.8, suppose that in addition $0<P[Y=i]<1$, $i=1,\ldots,d$, holds true. Then thanks to the fact that $\Omega_{Y}$ is finite, Assumptions 2.4 (i) and (ii) both hold true.
@@
$\frac{dP}{dP_{X}\otimes P_{Y}}\ =\ \varphi.$
@@
3 Factorizable joint shift (FJS) with general label spaces
He et al. (2021) proposed FJS “to handle the co-existence of sampling bias in covariates and labels”, and consequently as a notion generalising both covariate and label shift. Tasche (2022, Remark 1) characterised FJS in the case of categorical label variables as invariance up to constant factors of the ratios of the feature densities conditional on the labels between the source and the target distribution. In addition, he noted that in domain adaptation when only the feature marginal distribution of the target distribution is known (i.e. Problem (B) of Section 2), assuming FJS between the source and the target distributions does not uniquely determine the distribution shift. More recently, Dong et al. (2025) observed that FJS may be described as combination of a label and a covariate shift. This section generalises these observations for the case of FJS in the face of a general label space. In addition, solutions to Problems (A) and (B) of Section 2 under the assumption of FJS are discussed.
@@ 3.1 Characterising FJS
By definition (2.1), $\{X\in H,Y\in G:H\in\mathcal{H},G\in\mathcal{G}\}$ is a $\cap$-stable generator of $\mathcal{F}$, hence by the measure uniqueness theorem (Bauer, 1981, Theorem 1.5.5) $h(X)$ is a density of $Q$ with respect to $P$ on $(\Omega,\mathcal{F})$. Thus, $P$ and $Q$ are related through FJS with $\overline{h}=h$ and $\overline{g}=1$. Thanks to the symmetry of the setting with respect to $X$ and $Y$, we also have shown that label shift implies FJS with $\overline{h}=1$ and $\overline{g}=E_{P}[f\,|\,Y=\cdot]$. ∎
@@
    For all $F\in\mathcal{F}$ it holds that
    $\begin{split}Q_{L}[F\,|\,\sigma(Y)]\ =\ P[F\,|\,\sigma(Y)]\quad\text{and}\quad Q_{C}[F\,|\,\sigma(X)]\ =\ Q[F\,|\,\sigma(X)].\end{split}$ (3.2c)
    In other words, $P$ and $Q_{L}$ are related through label shift while $Q_{C}$ and $Q$ are related through covariate shift.
@@
    $E_{P}\bigl[\overline{g}(Y)\bigr]=0$ would imply $\overline{g}(Y)=0$ $P$-a.s. and therefore also $E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]=0$. This would contradict $E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]=1$ which is true by definition of FJS. This implies $E_{P}\bigl[\overline{g}(Y)\bigr]>0$. From the definition of $Q_{L}$ it follows that
@@
    Both equations in (iii) follow from Lemma 4.2 of Tasche (2024). ∎
@@
    Under Assumption 2.2, suppose that $h$ is a density of some probability measure $Q_{X}^{\ast}$ on $(\Omega_{X},\mathcal{H})$ with respect to $P_{X}$ and that $g$ is a density of some probability measure $Q_{Y}^{\ast}$ on $(\Omega_{Y},\mathcal{G})$ with respect to $P_{Y}$. In addition, let $\overline{h}:\Omega_{X}\to[0,\infty)$ be an $\mathcal{H}$-measurable function and $\overline{g}:\Omega_{Y}\to[0,\infty)$ be a $\mathcal{G}$-measurable function such that (3.3) holds true. Then $\overline{h}(X)\,\overline{g}(Y)$ is a probability density on $(\Omega,\mathcal{F})$, and for $Q$ defined by $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)$ it holds that $Q_{X}=Q_{X}^{\ast}$ and $Q_{Y}=Q_{Y}^{\ast}$.
@@
  • In general, $\overline{g}$ does not represent a marginal density of the label variable $Y$ as in the case of label shift but rather is characterised by being a conditional density of $Q_{Y|X}$ with respect to $P_{Y|X}$, after normalisation. Due to the symmetry of the setting in $X$ and $Y$, $\overline{h}$ is not a marginal density of the feature variable $X$ as in the case of covariate shift but rather is characterised by being a conditional density of $Q_{X|Y}$ with respect to $P_{X|Y}$, again after normalisation.
  • FJS is more homogeneous than non-specific distribution shift in so far as the conditional densities given the features are all equal (up to normalisation). ∎
3.2 Using FJS for inference about the target distribution
How to solve (3.3) for $\overline{h}$ and $\overline{g}$ if marginal densities $h$ and $g$ are given, i.e. if Problem (A) of Section 2 ought to be solved? Corollary 3.7 presents a transformation of (3.3) into an integral equation for $\overline{g}$ and a representation of $\overline{h}$ in terms of $h$ and $\overline{g}$.

### v2

@@ Abstract
Factorizable joint shift (FJS) represents a type of distribution shift (or dataset shift) that comprises both covariate and label shift. Recently, it has been observed that FJS actually arises from consecutive label and covariate (or vice versa) shifts. Research into FJS so far has been confined to the case of categorical labels. We propose a framework for analysing distribution shift in the case of a general label space, thus covering both classification and regression models. Based on the framework, we generalise existing results on FJS to general label spaces and present and analyse a related extension of the expectation maximisation (EM) algorithm for class prior probabilities. We also take a fresh look at generalized label shift (GLS) in the case of a general label space.
Keywords: Distribution shift, dataset shift, factorizable joint shift, generalized label shift, domain adaptation, EM algorithm.
@@ 1 Introduction
Classification and regression models may work very well on their respective training datasets but fail utterly when being deployed in test datasets. Such failure can be caused by distribution shift (also known as dataset shift) between the training and test datasets. For this reason, distribution shift and domain adaptation (a notion comprising techniques for tackling distribution shift) has been a major research topic in machine learning for some time. This paper takes the perspective of Kouw and Loog (2021) and studies the case where feature observations from the test dataset are available for analysis but observations of labels are missing.
Under these circumstances, without any assumptions on the nature of the distribution shift between the training and test datasets meaningful prediction of the labels in the test dataset or of their distribution is not feasible. See Kouw and Loog (2021) for a survey of approaches to domain adaptation and their related assumptions. Arguably, covariate shift (also known as population drift) and label shift (also known as prior probability shift or target shift) are the most popular specific distribution shift assumptions, both for their intuiveness as well as their computational manageability. However, exclusive covariate and label shift assumptions have been criticised for being insufficient for common domain adaptation tasks (e.g. Tachet des Combes et al., 2020).
He et al. (2021) proposed the notion of factorizable joint shift (FJS) as an assumption for distribution shift that covers both covariate and label shift. Recently, Dong et al. (2025) observed that FJS actually may be described as combination of label and covariate shifts, thus making the assumption more plausible. Tasche (2022) analysed FJS for categorical label spaces with regard to the potential and limitations of domain adaptation based on this assumption. This paper updates the analyses of Tasche (2022) for FJS in the case of a general label space. More precisely, the main contributions of this paper to the subject of distribution shift are the following:
  • Pointing out the building blocks of non-specific distribution shift with general label spaces (including Euclidean, categorical and mixed Euclidean-categorical label spaces) in order to facilitate the demarcation and analysis of specific types of distribution shift.
@@
  • Presenting a generalisation to general label spaces of the expectation maximisation (EM) algorithm for estimating target prior class probabilities by Saerens et al. (2002) and showing that it is fit for purpose.
@@
The setting and notation for this paper are set out in the first two sub-sections of Section 2, followed by a framework of general assumptions and results on non-specific distribution shift in the remaining sub-sections of Section 2. Section 3 comprises analyses and results regarding FJS as well as approaches to estimating the characterics of FJS, including the presentation and analysis of a general EM algorithm in Sub-section 3.2.3. The relation between GLS and FJS is analysed in Section 4. Section 5 concludes the paper, with a summary and suggestions for further research. Appendix A presents a description of the classification approach to density estimation while Appendix B contains the proofs for the results in Sub-section 3.2.3.
@@ 2 The building blocks of distribution shift
We study distribution shift related to classification and regression problems in a probabilistic setting with feature (also known as input, explanatory or independent) variables $X$ and label (also known as output, response or dependent) variables $Y$. We assume that a joint source (also known as training) probability distribution $P=P_{(X,Y)}$ of $X$ and $Y$ already has been learnt. Now we are presented with a joint target (also known as test) distribution $Q=Q_{(X,Y)}$ of $X$ and $Y$ which is possibly different to $P$ and only partially observable. Lack of observations from the target distribution $Q$ and more generally lack of information about $Q$ may become an issue in the presence of distribution shift (Zhang et al., 2024, Section 4.9). The problem then is to infer information about the posterior (also known as conditional) distribution $Q_{Y|X}$ of $Y$ given $X$ under the target distribution $Q$ from the source distribution $P$ despite a lack of joint observations of $(X,Y)$ under $Q$. In Section 3 below, we analyse the two cases where
@@
The problem phrased in (B) is an unsupervised domain adaptation problem where a classification or regression function is learnt on a completely known dataset in a supervised manner and is then adapted to another dataset without label information in an unsupervised manner (Kouw and Loog, 2021, Section 6.7). Problem (A) where only joint observations of features and label are missing has less frequently been looked at in the machine learning literature. It is akin to copula estimation (Durante and Sempi, 2016) and as such has some practical relevance (see e.g. Jaworski et al., 2013; Tasche, 2025). Problem (A) is investigated in this paper mainly because its solutions can inform approaches to solutions for Problem (B).
@@
  • knowledge of the building blocks can inform generative approaches to distribution shift when datasets need to be generated by simulation.
@@ 2.1 Setting
The setting for this paper is a standard setting for domain adaptation and therefore, with the exception of some details of the notation, similar to the setting described for example in the first paragraph of Section 2 of Azizzadenesheli (2022).
@@
  • The source distribution $P$ and the target distribution $Q$ are different, i.e. $P\neq Q$. In other words, there is a distribution shift (see e.g. Zhang et al., 2024) between source and target. Occasionally, non-specific distribution shift also is called dataset shift in the literature (Quiñonero-Candela et al., 2009; Moreno-Torres et al., 2012).
@@ 2.2 Notation
Expected values and probabilities conditional on $\sigma$-algebras or on random variables are well-defined for all combinations of probability spaces and integrable or non-negative random variables whose conditional expected values are meant to be determined (Bauer, 1981, Sections 10.1 and 10.2; Klenke, 2013, Sections 8.2 and 8.3).
@@
In general, the existence of regular conditional distributions can be guaranteed only if certain conditions on the image of the conditioned variable – $(\Omega_{Y},\mathcal{G})$ in Definition 2.1 – are satisfied (Bauer, 1981, Section 10.3; Klenke, 2013, Section 8.3). If however the label variable $Y$ is real-valued, the regular conditional distribution $P_{Y|X}$ exists, without any restrictions on the feature variable $X$ (Klenke, 2013, Theorem 8.28).
@@ 2.3 Linking source and target distributions in the face of distribution shift
In the literature the ‘transfer’ of information from $P$ to $Q$ by means of (2.5b) is sometimes called ‘importance reweighting’ and the density $f$ is referred to as ‘importance weights’ (e.g. Zhang et al., 2013; Kouw and Loog, 2021). More often than not, in the literature the reference measure for the density $f$ remains undisclosed or appears to be the Lebesgue measure (e.g. Zhang et al., 2013; Lipton et al., 2018). A notable exceptions are Scott (2019) who however proposes a reference measure $\mu$ that is available only in the case of a categorical label space, and Azizzadenesheli (2022).
@@
Since application of Fubini’s theorem (2.4a) cannot be justified in the absence of any assumptions on the properties of the source distribution $P$ in the setting as specified in Section 2.1, it is not possible to draw further conclusions about the composition of the density $f$ or the relationship between the posterior distributions of $X$ given $Y$ under the target distribution $Q$ and the source distribution $P$ respectively. This is in contrast to the situation studied by Tasche (2022). In Tasche (2022), thanks to the assumption that the label variable is categorical, more refined formulae for $f$ (Theorem 1) and the posterior probabilities (Corollary 2, a more general version of Eq. (2.4) of Saerens et al., 2002, but of similar shape) could be derived.
@@ 2.4 Assumptions on the source distribution and conditional densities
    Under Assumptions 2.2 and 2.4 (i), let $h$ be a density of $Q_{X}$ with respect to $P_{X}$ and define $q_{Y|X}:\Omega_{X}\times\Omega_{Y}\to[0,\infty)$ by $q_{Y|X=x}(y)=\begin{cases}\frac{f(x,y)}{h(x)},&h(x)>0\\ 0,&h(x)=0\end{cases}$ for $x\in\Omega_{X}$ and $y\in\Omega_{Y}$. Then for $x\in\Omega$ with $h(x)>0$ (and therefore $Q_{X}$-a.s.) $q_{Y|X=x}$ is a density of $Q_{Y|X=x}$ with respect to $P_{Y|X=x}$.
@@
    Under Assumptions 2.2 and 2.4 (ii), let $g$ be a density of $Q_{Y}$ with respect to $P_{Y}$ and define $q_{X|Y}:\Omega_{X}\times\Omega_{Y}\to[0,\infty)$ by $q_{X|Y=y}(x)=\begin{cases}\frac{f(x,y)}{g(y)},&g(y)>0\\ 0,&g(y)=0\end{cases}$ for $x\in\Omega_{X}$ and $y\in\Omega_{Y}$. Then for $y\in\Omega$ with $g(y)>0$ (and therefore $Q_{Y}$-a.s.) $q_{X|Y=y}$ is a density of $Q_{X|Y=y}$ with respect to $P_{X|Y=y}$.
@@
with a fixed arbitrary $y_{0}\in\Omega_{Y}$, for $x\in\Omega_{X}$ and $G\in\mathcal{G}$. Then $K$ satisfies (i) of Definition 2.1 and $K(x,\cdot)$ is a measure on $(\Omega_{Y},\mathcal{G})$. By Fubini’s theorem (2.4a), it follows for $H\in\mathcal{H}$ and $G\in\mathcal{G}$ that
@@
(i) is basically the definition of $q_{X|Y=y}(x)$ in Thorem 2.6 (ii). (ii) is the combination of (i) and Corollary 2.5 (i). (iii) is implied also by (i) and Corollary 2.5 (i). ∎
@@ 2.5 Facilitating calculations by further assumptions
Under Assumption 2.8, suppose in addition $0<P[Y=i]$, $i=1,\ldots,d$. Then thanks to the fact that $\Omega_{Y}$ is finite, Assumptions 2.4 (i) and (ii) both hold true.
@@
$\frac{dP}{dP_{X}\otimes P_{Y}}\ =\ \varphi.$ (2.8)
@@
3 Factorizable joint shift (FJS) for general label spaces
He et al. (2021) proposed FJS “to handle the co-existence of sampling bias in covariates and labels”, and consequently as a notion generalising both covariate and label shift. Tasche (2022, Remark 1) characterised FJS in the case of categorical label variables as invariance between the source and the target distribution up to constant factors of ratios of the feature densities conditional on different labels. In addition, he noted that in domain adaptation when only the feature marginal distribution of the target distribution is known (i.e. Problem (B) of Section 2), assuming FJS without constraints between the source and the target distributions does not uniquely determine the distribution shift. More recently, Dong et al. (2025) observed that FJS may be described as combination of a label and a covariate shift. This section generalises these observations for the case of FJS in the face of a general label space. In addition, solutions to Problems (A) and (B) of Section 2 under the assumption of FJS are discussed.
@@ 3.1 Characterising FJS
Conversely, if $P$ and $Q$ are related through FJS with $\overline{g}=1$ then $P$ and $Q$ are related also through covariate shift. If $P$ and $Q$ are related through FJS with $\overline{h}=1$ then $P$ and $Q$ are related also through label shift.
@@
By (2.1), $\bigl\{\{X\in H,Y\in G\}:H\in\mathcal{H},G\in\mathcal{G}\bigr\}$ is a $\cap$-stable generator of $\mathcal{F}$, hence by the measure uniqueness theorem (Bauer, 1981, Theorem 1.5.5) $h(X)$ is a density of $Q$ with respect to $P$ on $(\Omega,\mathcal{F})$. Thus, $P$ and $Q$ are related through FJS with $\overline{h}=h$ and $\overline{g}=1$. Thanks to the symmetry of the setting with respect to $X$ and $Y$, we also have shown that label shift implies FJS with $\overline{h}=1$ and $\overline{g}=E_{P}[f\,|\,Y=\cdot]$.
The converse claims follow from Lemma 4.2 of Tasche (2024). ∎
@@
    For all $F\in\mathcal{F}$ it holds that $Q_{L}[F\,|\,\sigma(Y)]=P[F\,|\,\sigma(Y)]$ and $Q_{C}[F\,|\,\sigma(X)]\ =\ Q[F\,|\,\sigma(X)]$. In other words, $P$ and $Q_{L}$ are related through label shift while $Q_{C}$ and $Q$ are related through covariate shift.
@@
    $E_{P}\bigl[\overline{g}(Y)\bigr]=0$ would imply $\overline{g}(Y)=0$ $P$-a.s. and therefore also $E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]=0$. This would contradict $E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]=1$ which is true by definition of FJS. Thus $E_{P}\bigl[\overline{g}(Y)\bigr]>0$ is true. From the definition of $Q_{L}$ it follows that
@@
    Both equations in (iii) follow from Proposition 3.3. ∎
@@
    Suppose that $h$ is a density of some probability measure $Q_{X}^{\ast}$ on $(\Omega_{X},\mathcal{H})$ with respect to $P_{X}$ and that $g$ is a density of some probability measure $Q_{Y}^{\ast}$ on $(\Omega_{Y},\mathcal{G})$ with respect to $P_{Y}$. In addition, let $\overline{h}:\Omega_{X}\to[0,\infty)$ be an $\mathcal{H}$-measurable function and $\overline{g}:\Omega_{Y}\to[0,\infty)$ be a $\mathcal{G}$-measurable function such that (3.3) holds true. Then $\overline{h}(X)\,\overline{g}(Y)$ is a probability density on $(\Omega,\mathcal{F})$, and for $Q$ defined by $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)$ it holds that $Q_{X}=Q_{X}^{\ast}$ and $Q_{Y}=Q_{Y}^{\ast}$.
@@
  • In general, $\overline{g}$ does not represent a marginal density of the label variable $Y$ as in the case of label shift (i.e. when $\overline{h}=1$) but rather is characterised by being a conditional density of $Q_{Y|X}$ with respect to $P_{Y|X}$, after normalisation. Due to the symmetry of the setting in $X$ and $Y$, $\overline{h}$ is not a marginal density of the feature variable $X$ as in the case of covariate shift but rather is characterised by being a conditional density of $Q_{X|Y}$ with respect to $P_{X|Y}$, again after normalisation.
  • FJS is more homogeneous than non-specific distribution shift in so far as the target conditional densities of the labels given the features are all equal (up to normalisation). ∎
3.2 Solving Problems (A) and (B) by assuming FJS
How to solve (3.3) for $\overline{h}$ and $\overline{g}$ if marginal densities $h$ and $g$ are given, i.e. if Problem (A) of Section 2 ought to be solved? Corollary 3.7 disentangles (3.3) to arrive at an integral equation for $\overline{g}$ and a representation of $\overline{h}$ in terms of $h$ and $\overline{g}$.